# Adaptive Reinforcement Learning Planning: Harnessing Large Language Models for Complex Information Extraction

### v1

@@ Abstract
Adaptive Reinforcement Learning Planning: Harnessing Large Language Models for Complex Information Extraction
Zepeng Ding, Ruiyang Ke, Wenhao Huang, Guochao Jiang, Yanda Li, Deqing Yang, Yanghua Xiao, Jiaqing Liang🖂 Shanghai Key Laboratory of Data Science School of Data Science, School of Computer Science, Fudan University {zpding22, ryke23, whhuang21, gcjiang22, ydli22}@m.fudan.edu.cn {yangdeqing, shawyh, liangjiaqing}@fudan.edu.cn
1 Introduction
@@
Current large language models (LLMs) have excellent performance in many fields such as dialogue, reading comprehension, and text generation, greatly expanding the boundaries of traditional NLP tasks and thus demonstrating the capacity to support complex information extraction (IE) tasks. Large models can understand extraction requirements through specific prompts and improve their performance on downstream tasks through in-context learning, multi-step planning Wei et al. (2023), collaborating with small model Ding et al. (2024) and so on. These methods achieve good performance in general IE tasks.
Figure 1: (a) Multi-step extraction reduces errors. (b) Extraction orders have an impact on the outputs.
@@
However, when confronted with complex situations such as very long sentences or paragraphs, multiple related entities in the same sentence, and tokenization confusion caused by overlapping tokens, LLMs still engender issues such as false positives and missing elements Ma et al. (2023); Wang et al. (2024), resulting relatively low F1-score. Fine-tuning LLMs with task data is a common solution to enhance extraction capabilities, but it is time-consuming, has limited generality and its performance highly relies on the quality of the labeled data. Even though LLMs are powerful, they still require multiple steps to achieve acceptable extraction results for complex cases Wei et al. (2023).
@@
Previous studies on small language models (SLMs) show that task decomposition and sequential extraction could reduce errors and achieve better results Xie et al. (2021); Lu et al. (2022). For LLMs, solving the task step by step rather than extracting all the content at once can deconstruct the information of the original sentence, leading to better extraction performance in complicated scenarios. An example is shown in Fig. 1. In our work, we divide the extraction process of LLMs into two distinct tasks: the classification of relations/events and the extraction of entities (arguments). We extract entities in multiple steps, orderly rather than concurrently. Indeed, many studies on LLM-based agents are using large models to solve problems in an orderly step-by-step manner with interaction and feedback, such as ReAct Yao et al. (2022) and AutoCrawler Huang et al. (2024). The effectiveness of these methods heavily depends on the inherent planning ability of LLMs.
@@
Taking complex information extraction as an example, many studies prove that it requires particularly high planning capabilities, which poses a considerable challenge to the aforementioned agents. First, the results of multi-step extraction are sensitive to the order of steps. Previous works use preset prompt templates and present a fixed order planning Wei et al. (2023); Ding et al. (2024). However, motivated by Huang et al. (2023), we find that varying extraction orders significantly influences the output of LLMs. An appropriate order can effectively utilize the logical association between entities and greatly reduce the complexity of the extraction task, thereby driving LLM to obtain better results. Second, the optimal order of entity extraction is not fixed for different sentences, even if the relation/event type is the same. As illustrated in Fig. 1, different extraction orders lead to different results, and even if the relation is the same as that in Fig. 1, the optimal order is distinct for these two sentences. It is difficult to obtain sufficient information to plan the optimal execution order for different sentences Li et al. (2024). Third, the optimal order for models of different capabilities and types is different, making planning and extraction execution entangled and difficult to decompose.
@@
Therefore, we believe that reinforcement learning is essential to generate plans that are acceptable to complex IE cases. In this paper, we establish a framework to effectively guide LLMs for multi-step extraction planning, which can have stable and effective extraction performance in different languages, sentence types, and tasks without fine-tuning. We propose a decision model to assist LLMs in executing entity extraction, adaptively deciding the subsequent action(s) based on the current state. Wed multi-step extractions as a Markov Decision Process (MDP), and design the state based on the input sentence, relation/event type, and extracted contents. We construct the environment with the LLM-based extractor, introduce an indicator as the reward function, and assign reward values based on the semantic correctness and token-level precision of the extraction results. After training through reinforcement learning and subsequently establishing a Deep Q-Network (DQN), the small decision model can collaboratively execute the sequential entity extraction tasks with the LLM-based extractor.
@@
  • First, we design a two-stage multi-step information extraction framework for LLMs, model order selection as MDP, and propose an easy-to-train decision model to guide LLMs in multi-step planning.
@@
2 Related Work
2.1 Research in Information Extraction
In Information Extraction (IE) tasks such as relation extraction and event extraction, research based on traditional models is relatively comprehensive. Pipeline IE methods dissect the extraction process into several sub-tasks, optimizing each individually. For instance, subjects, relations, and objects are sequentially extracted from sentences in pre-specified orders in addressing relation extraction tasks Wei et al. (2020); Xie et al. (2021). When dealing with event extraction, event types, and triggers are initially identified, followed by the extraction of the arguments using sequence tagging models or machine comprehension models Yang et al. (2019). Joint IE methods combine several extraction processes into a single stage. For example, graph-based approaches identify the entity or text span and construct a graph with relations or co-occurrences Wadden et al. (2019); Zhu et al. (2021), while generation-based approaches serialize structured extraction results into sentences or pre-defined templates in a fixed order Hsu et al. (2022); Ye et al. (2021); Lu et al. (2022).
@@ 2.1 Research in Information Extraction
Recent studies on Large Language Models (LLMs) demonstrate their proficient performance across various downstream tasks, even when provided with only a few examples as prompts Agrawal et al. (2022); Jeblick et al. (2023). For IE tasks, some studies indicate that with appropriate prompting, LLMs can achieve comparable performance with the supervised methods in zero-shot or few-shot settings of extraction tasks Wei et al. (2023); Gao et al. (2023); Tang et al. (2023). Ding et al. (2024) integrates LLMs with a small evaluation model for RE tasks, addresses the "low recall" issue, and enhances the extraction performance of LLMs. Furthermore, Chung et al. (2022) shows LLMs can produce outstanding performance through supervision and fine-tuning, and Wadhwa et al. (2023) suggests that LLMs should be a standard benchmark for some extraction tasks.
2.2 Reinforcement Learning in IE Tasks
@@
Previous studies have also employed reinforcement learning (RL) in IE tasks. Narasimhan et al. (2016) enhances information extraction by utilizing RL to integrate external evidence. Takanobu et al. (2019) implements a hierarchical RL framework to enhance the linkage between entity mentions and relation types. Zeng et al. (2018) uses the relation of entity pairs as distant supervision and guides the training of the relation extractor with RL to address large-scale RE tasks. The ordered extraction of entities can be modeled as a sequential decision-making problem, so it can be naturally integrated with RL. For example, Zeng et al. (2019) contemplates the extraction order of relational facts in a sentence and subsequently trains a sequence-to-sequence model using RL, and Huang et al. (2023) proposes an RL-based framework that dynamically generates an optimal extraction order for each sentence to guide the model towards better extraction performance.
3 Methods
3.1 Preliminary and Task Definition
@@
Information extraction (IE) tasks are highly diversified due to their different targets Lu et al. (2022). For this paper, we mainly focus on relation extraction (RE) and event extraction (EE) tasks. RE task aims at extracting triples (subject, predicate, object) (or (s,p,o)) from a given natural language sentence. The subject entity and object entity are linked by a specific predicate (relation). EE task intends to identify the event types from the given sentence and identify the arguments (entities) corresponding to each role. The extracted arguments need to be consistent with the event type and semantically and logically correct.
Figure 2: An example of the multi-step information extraction of EE task.
We uniformly define RE and EE as a two-stage task, as shown in Fig.  2. The goal of the first stage is to classify the relation or event type based on the given sentence, and the second stage is to extract the corresponding entities or arguments. More specifically, the second stage can be formulated as a multi-argument extraction process: given the sentence $s$, the relation/event type $pred$, our goal is to extract all the arguments in $s$ and correctly fill them into the schema $[role_{1},role_{2},...,role_{n}|pred]$. In RE tasks, $n=2$ for subject and object.
Figure 3: An example of modeling the multi-step entity extraction as MDP. In each step, the decision module chooses a role, and the LLM-based environment returns the corresponding argument(s) and reward(s).
3.2 Solution Framework
@@
Generally, for the multi-step argument extraction task, the LLM-based extractor acts as an environment to provide the decision model with the state after the current extraction execution, and the decision model selects the most appropriate role based on the state to guide LLMs to extract step by step until all roles are extracted.
@@
We use the LLMs in both stages without fine-tuning with labeled data. The output of LLMs is crucial to completing the RL training of the decision model. To make the output of LLMs more stable, and to enable the LLM-based environment to effectively interact with the decision module, we model the reward assignment as a classification task and design a reward model that takes into account both semantic correctness and token-level matching. The reward model and extraction prompts are described in subsequent subsections.
For the classification task, we provide LLMs with input text and a list of candidate relation/event types and also provide several appropriate examples for different datasets in the prompt to guide LLMs to select the type(s) from the candidate list  [^1] LLMs can easily follow instructions to make correct classifications (also verified in Table 4).. For argument extraction in the second stage, each step only requires LLMs to extract arguments (entities) corresponding to a single role. In the $i$-th step, we provide the input text $s$ and the previously extracted role-argument pairs $(r_{j}:a_{j})_{j=1:i-1}$ as input to LLMs, and instruct LLMs to extract the argument $a_{i}$ corresponding to the currently required role $r_{i}$.
3.3 MDP for Argument Extraction
@@
We build the environment with the LLM-based extractor, initialize the state based on the relation/event extracted in the first stage, and determine the roles schema (as the initial action space). In each step, the decision model takes the input sentence and extracted content as input and chooses a previously unselected role as the action. The environment receives the action from the decision model, constructs the extractor’s input, and performs extraction. After collecting the extraction results, the environment would assign the reward based on the ground truth of the labeled sample, separate the different arguments of the current role, build the successor state(s), and transit to a new state. When all roles are selected (the action space is empty), the decision model returns a termination signal, and the environment converts the extracted content into structural output. Fig. 3 shows the entire extraction process of an instance.
@@
Where the extracted content consists of previously extracted role-argument pairs: $C_{i}=(r_{j}:a_{j})_{j=1:i-1}$. In each step, the environment constructs the input of LLM-based extractor according to the current state $S_{i}$ and the role $r_{i}$ selected by the decision model and transits to a new state.
@@
For a given sentence and corresponding relation/event type, the action of the decision model is the next role to extract. The initial action space $A_{0}$ is the roles schema $\mathcal{M}$ in the state. In a complete argument extraction process, the extraction order is a permutation of roles in $\mathcal{M}$, and each role should be extracted only once. Therefore, the action space of the decision model is reduced at each step. After selecting action (role) $a_{i}=r_{i}$ at the $i$-th step, $a_{i}$ will be removed from the action space $A_{i}$ to derive $A_{i+1}$. The update of the action space can be expressed by the following formula:
$A_{i}=\begin{cases}\mathcal{M},&i=0\\ A_{i-1}-\{a_{i-1}\},&0<i<|\mathcal{M}|\end{cases}$ (2)
Reward
@@
The design of rewards is crucial in RL training and is directly related to the loss and the update of model weights. In our framework, the reward value is used to measure the quality of the extraction results and guide the decision model to make better plans for each episode. However, the output of a large model is usually uncertain and imprecise. On one hand, applying the exact match with the ground-truth will misjudge many semantically correct answers as wrong, reducing the effectiveness of the reward value. On the other hand, IE tasks require accurate results in some specific areas and scenarios. Therefore, we hope to assign high reward values to outputs that are both semantically correct and highly matched with the ground-truth, and use low reward values to guide LLMs to avoid outputs that are semantically incorrect or mismatched with the ground-truth. Details of our reward function and reward model are as follows.
3.3.1 Reward Function and Reward Model
@@
Based on the above considerations, we set the reward function to an indicator function, that is, to assign a reward value of 1 to acceptable extraction results [^2] To match the magnitude of the $Q$ value generated by the decision model, in practical we use 0 or 10 as the reward., and to assign a reward value of 0 to semantically incorrect or mismatched answers. We directly use an LLM to evaluate the semantic correctness, automatically implement this concise but effective reward function, and use token-level similarity between the extracted results and the ground-truth as a threshold to help determine whether they match.
@@
Note that our reward function has only two discrete values (which can be regarded as two labels), rather than a continuous function. On one hand, it is intuitive and reasonable to classify the extraction results into "acceptable" and "need to be rejected", but it is difficult to find a reliable and stable method to further sort the acceptable results and assign different rewards. On the other hand, using an LLM as our reward model is convenient and effective. It only needs to judge whether the extraction results are acceptable and output 0 or 1.
@@
During the RL training process, we use an LLM with a larger scale and stronger capabilities than the extraction model as the reward model to ensure accurate and effective reward assignment. This process also performs implicit knowledge distillation through the rewards. We instruct it to consider whether the semantics of the extracted arguments are consistent with the ground-truth and the token-level accuracy. See Appendix C for the prompt template.
@@
We use the word-level textF1 score Huang et al. (2023) defined as follows to measure the token-level similarity between ground-truth and extracted arguments and use it as a threshold to remove "mismatched" extraction results. This step can be regarded as a pruning process, which reduces the successor state space and enhances the effectiveness when performing inference.
$Set_{a}=$ $Set_{p}\cap Set_{g}$ (3)
$Text\text{P}=\frac{|Set_{a}|}{|Set_{p}|},$ $\quad Text\text{R}=\frac{|Set_{a}|}{|Set_{g}|}$
[…]
3.4 Training of Decision Model
@@
We use a BERT-based decision model to evaluate the value of (State, Action) pair to provide the next action for the environment. Specifically, we use BERT Devlin et al. (2019) for English datasets and RoBERTa Liu et al. (2019) for Chinese datasets. We concatenate the input state and action into strings to form a sequence $x_{t}$ = [[CLS],$\mathbf{a_{t}}$,[SEP],$\mathbf{S_{t}}$,[SEP]].
@@
The Bert encoder converts these tokens into hidden vectors, and we select $h_{0}$ (the encoding vector of the token [CLS]) as the representative of the (State, Action) pair. The final evaluation of $Q(S_{t},a_{t})$ is defined as:
$\hat{Q}=\mathbf{Wh_{0}}+\mathbf{b}$ (4)
@@
Input: $\mathcal{D}:$initialized replay memory; $\theta$: network parameters; $\theta^{\prime}$: target-net parameters; $N_{b}$:training batch size
Input: hyper-parameters $E,\epsilon,\gamma,k$
1for epoch=1,…,$E$ do
2       Sample instances from labeled train set.
3       Initialize the state $S_{0}=(\mathcal{M}_{0},C_{0},p,s)$.
4       for $|\mathcal{M}|$ do
5             $p=$ Random$(0,1)$
6             if p<1-$\epsilon$ then
7                   $a_{t}=\arg\max_{a}Q(S_{t},a;\theta)$
8            else
9                   $a_{t}=$ Random-sample$(\mathcal{A}_{t})$
10             end if
11            $S_{t+1},r_{t}=$ Transit$(s_{t},a_{t})$
12             Store $(S_{t},a_{t},r_{t},S_{t+1})$ in $\mathcal{D}$
13             Sample random mini-batch of $N_{b}$ transitions $(S_{j},a_{j},r_{j},S_{j+1})$ from $\mathcal{D}$, and do follows:
14             if $S_{j+1}$ is terminal state then
15                   $y_{j}=r_{j}$
16            else
17                   $y_{j}=r_{j}+\gamma\max_{a}Q(S_{j+1},a;\theta^{\prime})$
18             end if
19            Update parameter $\theta$ on the loss $\mathcal{L}(\theta)=(y_{j}-Q(S_{t},a_{t};\theta))^{2}$
20             Replace target-net parameters $\theta^{\prime}$ = $\theta$ every $k$ steps
21       end for
22
23 end for
Algorithm 1 Training of decision model: DDQN algorithm with experience replay
We apply Deep Q-Learning to train the model and build a Deep Q-Network (DQN). Unlike the traditional DQN Mnih et al. (2013), in our tasks, a role may correspond to multiple arguments, and the extractor could extract multiple answers (arguments) based on the current state, or may not return any results (if there is no argument that matches the extracted content and the role). We design the following learning object:
@@ 20             Replace target-net parameters $\theta^{\prime}$ = $\theta$ every $k$ steps
We adopt the DDQN algorithm Van Hasselt et al. (2016), utilize $\epsilon-$greedy exploration and the Experience Replay Mnih et al. (2013) for RL training, as illustrated in Algorithm 1. The loss function is defined as the expected value of the mean squared Temporal Difference (TD) error:

### v2

@@ Abstract
Introduction
@@
Current large language models (LLMs) have excellent performance in many fields such as dialogue, reading comprehension, and text generation, greatly expanding the boundaries of traditional NLP tasks and thus demonstrating the capacity to support complex information extraction (IE) tasks. Large models can understand extraction requirements through specific prompts and improve their performance on downstream tasks through in-context learning, multi-step planning (Wei et al. 2023), collaborating with small model (Ding et al. 2024) and so on. These methods achieve good performance in general IE tasks.
Figure 1: (a) Extraction orders influence the outputs. (b) Planning and extracting are entangled for each model.
@@
However, when confronted with complex situations such as very long sentences or paragraphs, multiple related entities in the same sentence, and tokenization confusion caused by overlapping tokens, LLMs still engender issues such as false positives and missing elements (Ma et al. 2023; Wang et al. 2024), resulting relatively low F1-score. Fine-tuning LLMs with task data is a common solution to enhance extraction capabilities, but it is time-consuming, has limited generality and its performance highly relies on the quality of the labeled data. Even though LLMs are powerful, they still require multiple steps to achieve acceptable extraction results for complex cases (Wei et al. 2023).
@@
Previous studies on small language models (SLMs) show that task decomposition and sequential extraction could reduce errors and achieve better results (Xie et al. 2021; Lu et al. 2022). In our work, we divide the extraction process of LLMs into two distinct tasks: the classification of relations/events and the extraction of entities (arguments). We extract entities in multiple steps, orderly rather than concurrently. Indeed, many studies on LLM-based agents are using large models to solve problems in an orderly step-by-step manner with interaction and feedback, such as ReAct (Yao et al. 2022) and AutoCrawler (Huang et al. 2024). The effectiveness of these methods heavily depends on the inherent planning ability of LLMs.
@@
Taking complex information extraction as an example, many studies prove that it requires particularly high planning capabilities, which poses a considerable challenge to the aforementioned agents. First, the results of multi-step extraction are sensitive to the order of steps. Huang et al. (2023) verifies that the extraction order affects the results of the BERT-based model, but the structure of the generative large model and its performance on language tasks differ from those of the traditional small model, so the extraction performance of LLMs deserves further study [^1] In IE tasks, traditional models tend to extract excessive entities, but LLMs tend to get a low recall result (Ding et al. 2024).. Previous works on LLMs use preset prompt templates and present a fixed order planning (Wei et al. 2023; Ding et al. 2024). However, we find that varying extraction orders significantly influences the output of LLMs, and the optimal order of entity extraction is not fixed for different sentences. As illustrated in Fig. 1, different extraction orders lead to different results, and the optimal order is distinct from the situation in Fig. 1. Second, the optimal order for models of different capabilities and types is different, making planning and extraction execution entangled and difficult to decompose. If the previous planning causes extraction errors in the intermediate results, the wrong results will also affect the execution of subsequent step(s). As illustrated in Fig. 1, LLM A can identify the song “Hotel California” based on its inherent common sense knowledge, so both plans can get the correct results. For LLM B, the plan “find the object first” can guide it to extract “Hotel California” based on the sentence structure, but if our plan is “find the subject first”, the model tends to extract “California” as the location and affect the extraction of object. Consequently, it is difficult to obtain sufficient information to plan the optimal execution order by relying solely on the LLMs’ planning ability.
@@
Therefore, we believe that reinforcement learning is essential to generate plans that are acceptable to complex IE cases. In this paper, we establish a framework to effectively guide LLMs for multi-step extraction planning, which can have stable and effective extraction performance in different languages, sentence types, and tasks without fine-tuning. We propose a decision model to assist LLMs in executing entity extraction, adaptively deciding the subsequent action(s) based on the current state. We construct the environment with the LLM-based extractor, introduce an indicator as the reward function, and assign reward values based on the semantic correctness and token-level precision of the extraction results. After training through reinforcement learning and subsequently establishing a Deep Q-Network (DQN), the BERT-based decision model can collaboratively execute the sequential entity extraction tasks with the LLM-based extractor. In other words, we adopt model collaboration to deal with the entanglement of planning and extracting, let the decision model and the LLM “do what they are good at respectively” (LLM for language understanding and extracting, and the decision model for fitting Q values and planning), and combine them through RL.
@@
  • First, we design a two-stage multi-step information extraction framework for LLMs, model order selection as Markov Decision Process (MDP), and propose an easy-to-train decision model to guide LLMs in multi-step planning.
@@
Related Work
LLM for Information Extraction
@@
Recent studies on LLMs demonstrate their proficient performance across various downstream tasks, even when provided with only a few examples as prompts (Agrawal et al. 2022; Jeblick et al. 2023). For IE tasks, some studies indicate that with appropriate prompting, LLMs can achieve comparable performance with the supervised methods in zero-shot or few-shot settings of extraction tasks (Wei et al. 2023; Gao et al. 2023; Tang et al. 2023). Ding et al. (2024) integrates LLMs with a small evaluation model for RE tasks, addresses the ”low recall” issue, and enhances the extraction performance of LLMs. Furthermore, Chung et al. (2022) shows LLMs can produce outstanding performance through supervision and fine-tuning, and Wadhwa, Amir, and Wallace (2023) suggests that LLMs should be a standard benchmark for some extraction tasks.
Reinforcement Learning in IE Tasks
@@
Previous studies have also employed reinforcement learning (RL) in IE tasks. Narasimhan, Yala, and Barzilay (2016) enhances information extraction by utilizing RL to integrate external evidence. Takanobu et al. (2019) implements a hierarchical RL framework to enhance the linkage between entity mentions and relation types. Zeng et al. (2018) uses the relation of entity pairs as distant supervision and guides the training of the relation extractor with RL to address large-scale RE tasks. The ordered extraction of entities can be modeled as a sequential decision-making problem, so it can be naturally integrated with RL. For example, Zeng et al. (2019) contemplates the extraction order of relational facts in a sentence and subsequently trains a sequence-to-sequence model using RL, and Huang et al. (2023) proposes an RL-based framework that simultaneously trains BERT-based extraction and decision models to generate a suitable extraction order under a given schema.
Methods
Figure 2: The main workflow of our method. On the left is the extracting process, including relation/event classification and arguments extraction. On the right is the planning part, which guides each step based on the BERT-based Q-Network.
Preliminary and Task Definition
@@ Methods
Information extraction (IE) tasks are highly diversified due to their different targets (Lu et al. 2022). For this paper, we mainly focus on relation extraction (RE) and event extraction (EE) tasks. RE task aims at extracting triples (subject, predicate, object) (or (s,p,o)) from a given natural language sentence. EE task intends to identify the event types from the given sentence and identify the arguments (entities) corresponding to each role. We uniformly define RE and EE as a two-stage task. The goal of the first stage is to classify the relation or event type based on the given sentence, and the second stage is to extract the corresponding entities or arguments. The extracted arguments need to be consistent with the relation/event type and semantically and logically correct.
For the relation/event classification, we provide LLMs with input text and a list of candidate relation/event types and also provide several appropriate examples for different datasets in the prompt to guide LLMs to select the type(s) from the candidate list. For argument extraction in the second stage, each step only requires LLMs to extract arguments (entities) corresponding to a single role. In the $i$-th step, we provide the input text $s$ and the previously extracted role-argument pairs $(r_{j}:a_{j})_{j=1:i-1}$ as input to LLMs, and instruct LLMs to extract the argument $a_{i}$ corresponding to the currently required role $r_{i}$.
Solution Framework
@@
In previous work, Huang et al. (2023) models BERT-based entity extraction as MDP. However, their decision model and extraction model have the same structure and shared logits, and they are co-trained based on the same training data. Since the fine-tuning of extraction model can also improve the capability, it is not certain whether the final performance improvement comes from the appropriate planning.
@@
In our study, we decouple planning and extracting processes, as shown in Fig. 2. Generally, we use the LLMs in both stages without fine-tuning with labeled data. For the multi-step argument extraction task, the LLM-based extractor acts as an environment to provide the decision model with the state after the current extraction execution, and the decision model selects the most appropriate role based on the state to guide LLMs to extract step by step until all roles are extracted. The training of the decision model is separated from the execution of the whole extraction process, and the trained decision module can directly interact with similar extraction environments and serve as an instant, universal planner.
@@
The output of LLMs is crucial to completing the RL training of the decision model. To make the output of LLMs more stable, and to enable the LLM-based environment to effectively interact with the decision module, we model the reward assignment as a classification task and design a reward model that takes into account both semantic correctness and token-level matching. The reward model and extraction prompts are described in subsequent subsections.
LLM for Relation/Event Classification
In addition, in some datasets, the relation/event types of the originally labeled data are not natural language expressions. Consequently, it is challenging for LLMs to comprehend the semantics of these relation/event words without undergoing fine-tuning training. To address this, we undertake a natural language conversion for these labels, such as converting ”people-deceased_person-place_of_death” in NYT10 dataset to ”dead in”. Please refer to Technical Appendix for prompt template.
LLM-Based Argument Extractor
We adopt an LLM-based extractor to solve the argument extraction task with different extraction orders. For EE tasks, the roles are determined by the event type, and the event type has a corresponding roles schema. The extractor extracts the arguments for the selected role. For RE tasks, the intuitive approach is to set two roles for all relation types: subject and object. However, unlike roles in EE, the words ”subject” and ”object” do not contain any information about the entity type, which sometimes leads to ambiguity. For example, the subject-object of the relation ”affiliated” is person-organization in some datasets, while in other datasets it is department-organization. Therefore, we design the roles in RE tasks as ”subject/object: entity type”. The entity type is determined by the semantics of the selected relation type in the specific dataset.
The input of the arguments extractor is a dictionary, including the original sentence, the current relation/event type, the extracted content (role-argument pairs), and the role (argument type) to extract. At each step of argument extraction, the extractor returns the arguments corresponding to the specified role in a string, separated by commas. The prompt template is provided in Technical Appendix.
Environment for Extraction
@@
We build the environment with the LLM-based extractor, initialize the state based on the relation/event extracted in the first stage, and determine the roles schema (as the initial action space). In each step, the decision model takes the input sentence and extracted content as input and chooses a previously unselected role as the action. The environment receives the action from the decision model, constructs the extractor’s input, and performs extraction. After collecting the extraction results, the reward module would assign the reward based on the ground truth of the labeled sample, separate the different arguments of the current role, build the successor state(s), and transit to a new state. When all roles are selected (the action space is empty), the decision model returns a termination signal, and the environment converts the extracted content into structural output. The left of Fig. 2 shows the composition of the environment in detail.
@@
Where the extracted content consists of previously extracted role-argument pairs: $C_{i}=(r_{j}:a_{j})_{j=1:i-1}$.
@@
For a given sentence and corresponding relation/event type, the action of the decision model is the next role to extract. The initial action space $A_{0}$ is the roles schema $\mathcal{M}$ in the state. In a complete argument extraction process, the extraction order is a permutation of roles in $\mathcal{M}$, and each role should be extracted only once. Therefore, the action space of the decision model is reduced at each step. After selecting action (role) $a_{i}=r_{i}$ at the $i$-th step, $a_{i}$ will be removed from the action space $A_{i}$ to derive $A_{i+1}$.
The Reward Module
@@
The design of rewards is crucial in RL training and is directly related to the loss and the update of model weights. In our framework, the reward value is used to measure the quality of the extraction results and guide the decision model to make better plans for each episode. However, the output of a large model is usually uncertain and imprecise. On one hand, applying the exact match with the ground-truth will misjudge many semantically correct answers as wrong, reducing the effectiveness of the reward value. On the other hand, IE tasks require accurate results in some specific areas and scenarios. Therefore, we hope to assign high reward values to outputs that are both semantically correct and highly matched with the ground-truth, and use low reward values to guide LLMs to avoid outputs that are semantically incorrect or mismatched with the ground-truth. We design an indicator function to judge the semantic correctness and apply text similarity to remove results with low accuracy.
@@
We set the reward function to an indicator function, that is, to assign a reward value of 1 to acceptable extraction results [^2] To match the magnitude of the $Q$ value generated by the decision model, in practice we use 0 or 10 as the reward., and to assign a reward value of 0 to semantically incorrect or mismatched answers. We directly use an LLM to evaluate the semantic correctness, automatically implement this concise but effective reward function.
@@
Note that our reward function has only two discrete values (which can be regarded as two labels), rather than a continuous function. On one hand, it is intuitive and reasonable to classify the extraction results into ”acceptable” and ”rejective”, but it is difficult to find a reliable and stable method to further sort the acceptable results and assign different rewards. On the other hand, using an LLM as our reward model is convenient and effective. It only needs to judge whether the extraction results are acceptable and output 0 or 1.
Algorithm 1 Training of decision model
Input: Train sets $\mathcal{T}_{1},...,\mathcal{T}_{T}$ with the same task type and language; Initialized replay memory $\mathcal{D}$
Parameter: $\theta$: network parameters; $\theta^{\prime}$: target-net parameters; $N_{b}$:training batch size; $T_{h}$: Threshold of text-similarity; hyper-parameters $E,\epsilon,\gamma,k$
1:  for epoch$=$1,…,$E$ do
2:     Sample instance $s$ from $T$ labeled train sets, with probability $1/T$ for each set.
3:     Get the relation or event type $p$, and corresponding roles schema $\mathcal{M}_{0}$.
4:     Initialize the state $S_{0}=(\mathcal{M}_{0},C_{0},p,s)$.
5:     for $|\mathcal{M}|$ do
6:        $p=$ Random$(0,1)$
7:        if p $\leq$ 1-$\epsilon$ then
8:           $a_{t}=\arg\max_{a}Q(S_{t},a;\theta)$
9:        else
10:           $a_{t}=$ Random-sample$(\mathcal{A}_{t})$.
11:        end if
12:        $aug_{t+1}=$ Extractor$(S_{t},a_{t})$
13:        Get ground truth $g_{t+1}$ from instance $s$.
14:        if text-similar$(aug_{t+1},g_{t+1})\leq T_{h}$ then
15:           $r_{t+1}=0$
16:        else
17:           $r_{t+1}=$ Reward-Model$(aug_{t+1},g_{t+1})$
18:        end if
19:        Construct $S_{t+1}$, and store $(S_{t},a_{t},r_{t+1},S_{t+1})$ in $\mathcal{D}$
20:        Randomly sample $N_{b}$ transitions $(S_{j},a_{j},r_{j},S_{j+1})$ from $\mathcal{D}$, and do follows:
21:        if $S_{j+1}$ is terminal state then
22:           $y_{j}=r_{j}$
23:        else
24:           $y_{j}=r_{j}+\gamma\max_{a}Q(S_{j+1},a;\theta^{\prime})$
25:        end if
26:        $\mathcal{L}(\theta)=(y_{j}-Q(S_{t},a_{t};\theta))^{2}$
27:        Update parameter $\theta$
28:        Replace target-net parameters $\theta^{\prime}$ = $\theta$ every $k$ steps
29:     end for
30:  end for
@@
During the RL training process, we use an LLM with a larger scale and stronger capabilities than the extraction model as the reward model to ensure accurate and effective reward assignment. This process also performs implicit knowledge distillation through the rewards. We instruct it to consider whether the semantics of the extracted arguments are consistent with the ground-truth and the token-level accuracy. See Technical Appendix for the prompt template.
@@
We use the token-level text similarity (Jaccard Similarity) between ground-truth and extracted argument to measure the token-level accuracy and use it as a threshold to remove ”mismatched” extraction results. After LLM extraction, we first calculate the similarity score of each argument based on the ground-truth, and set the reward corresponding to the low-score arguments to 0 without passing them to the reward function. This step can be regarded as a pruning process, which reduces the successor state space and enhances the effectiveness when performing inference. The entire reward module can be formulated as follows:
$R(x)=\begin{cases}1,&\text{if}\ Sim_{x,g}>T\ \text{and}\ Ind_{x,g}=1\\ 0,&\text{otherwise.}\end{cases}$ (2)
Where $g$ is the ground-truth, $Sim_{x,g}$ refers to the text similarity between $x$ and $g$.
Decision Module for Planning
@@
We use a BERT-based decision model to evaluate the value of (State, Action) pair to plan the next action for the environment. At each step, we combine each role $a_{t_{i}}$ in the action space with the state $S_{t}$ as the input sequence: $x_{t_{i}}$ = [[CLS],$a_{t_{i}}$,[SEP],$S_{t}$,[SEP]].
@@
The BERT encoder converts these tokens into hidden vectors, and we select the first vector $h_{0}$ as the representative of the (State, Action) pair. The final evaluation of $Q(S_{t},a_{t})$ is defined as:
$\hat{Q}(S_{t},a_{t})=\mathbf{W}h_{0}+\mathbf{b}$ (3)
@@
We apply Deep Q-Learning to train the model and build a Deep Q-Network (DQN) (Mnih et al. 2013). We design the following learning object:
@@
We adopt the DDQN algorithm (Van Hasselt, Guez, and Silver 2016), utilize $\epsilon-$greedy exploration and the Experience Replay (Mnih et al. 2013) for RL training. The loss function is defined as the expected value of the mean squared Temporal Difference (TD) error: